OpenStudy (anonymous):

If X is a binomial random variable having density function binom(n,p) , we approximate P(X≤k) by P(Znp,np(1−p)≤k+1/2) . Use this approximation in the following. If X has binomial density function binom(6000,0.6) , approximate P(X≤3620) : P(X≤3620)≐P(Za,b≤c)

OpenStudy (anonymous):

what kind of table or TI or maple commands i need to find this: This probability is thus P(Za,b≤c)=? I know that A= 3600 np b=37.9473 \[\sqrt{np*(1-p)}\] and finally c= 3620.5 \[3620+\frac{ 1 }{ 2 }\]